4x squared equals 35 minus 4x
Answer by stanbon(57347)

4x squared equals 35 minus 4x
------------------
4x^2 = 35 - 4x
---
Rearrange:
4x^2 + 4x - 35 = 0
(2x+7)(2x-5) = 0
x = -7/2 ; x = 5/2
=====================
Cheers,
Stan H.